### v1

# Know the Unknown: An Uncertainty-Sensitive Method for LLM Instruction Tuning

Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across various tasks but still face challenges such as hallucinations. One potential reason for hallucinations is the lack of relevant knowledge or context. Thus, a promising solution to mitigate this issue involves instructing LLMs to respond with "I do not know" when a question falls outside their knowledge domain or the provided context. However, in this work, we observed that LLMs struggle to admit their lack of knowledge, primarily due to existing instruction datasets designed to encourage specific answers. To improve large language models’ capability to recognize the boundaries of their knowledge, we propose a novel approach called uncertainty-sensitive tuning. This method involves two-stage training designed for uncertainty recognition and prompt-sensitive activation. In the first stage, we guide the LLM to reject unknown questions. In the second stage, we recover the decreased performance in QA tasks by incorporating designed causal instructions. By leveraging this method, we aim to enhance the model’s ability to identify areas of uncertainty. The experimental results demonstrate that our proposed uncertainty-sensitive tuning method significantly improves the performance of the Llama2-chat-7B model. Specifically, it achieves a substantial 34.7% improvement in handling questions involving knowledge gaps compared to the original model. Moreover, our approach outperforms GPT-4, exhibiting a 9.4% increase in overall performance. We open-source the model and code on GitHub [^1] https://github.com/JiaqiLi404/TrustworthyRAG.
1 Introduction
The advancement of large language models (LLMs) has shown great potential for various tasks Brown et al. (2020); Wei et al. (2022); Joshi et al. (2017). However, there are still some issues, such as hallucinations Huang et al. (2023); Ji et al. (2023); Ye et al. (2023). Hallucination is a phenomenon in which LLMs generate a response that is either factually incorrect or nonsensical Maynez et al. (2020). Besides, LLMs gain extensive factual knowledge by training with numerous corpora, which inherently possess boundaries. When faced with unfamiliar information, the LLM’s response tends to be nonsense or outdated Huang et al. (2023).
Figure 1: The response of different models facing two types of questions. Known Question: The question has specific answers. Unknown Question: The question falls outside the provided context.
To bridge the knowledge gap, there are two main approaches. The first approach is to add new knowledge to the model, such as further fine-tuning Liu et al. (2023b); Gao et al. (2023); Liu et al. (2023a). The second approach is to equip the LLMs with an external database and use retrieval-augmented generation Es et al. (2023). However, the existing instruction dataset focuses on teaching LLMs to answer questions with a specific answer, making models incapable of responding "I don’t know" Zhang et al. (2023a), which will also increase the tendency for hallucination. To address this issue, many researchers Zhou et al. (2023); Yu et al. (2024) suggest prompting LLMs to admit their lack of knowledge. However, by doing this, LLMs may ignore important instructions contained in the prompts Kasai et al. (2023); Li et al. (2023); Si et al. (2023). As shown in Fig. 1, LLMs may still utilize knowledge outside the provided corpus even with explicit instructions, such as "If the context is not sufficient to answer the question, please answer it with ’Not Provided’.".
This paper is motivated by the findings that when there are tiny discrepancies between the knowledge and the question, or the generation requires multi-hop reasoning Huang et al. (2023); Shuster et al. (2021), LLMs tend to be hallucinating. Thus, we focus on the situation in which the provided context is insufficient. Initially, we explore why LLMs struggle to acknowledge their limitations in knowledge. We hypothesize that this issue derives from the models’ perceptual capabilities and introduces uncertainty-aware assessment. We utilize a dataset featuring known and unknown questions, where the unknown questions are crafted with insufficient context for answering questions. These samples are generated using the ASQA dataset Stelmakh et al. (2022), which contains 3,320 ambiguous question-context pairs. We pair each question with a slightly insufficient context. Compared to other works Li et al. (2022); Chen et al. (2023), the slight difference could significantly increase the level of model uncertainty. This dataset, called the uncertainty-recognition benchmark, serves as the foundation for our performance assessments.
To enhance the capabilities of large language models (LLMs) in recognizing the limits of their knowledge, we introduce a novel training framework called uncertainty-sensitive tuning. This approach contains a two-stage training process specifically tailored for uncertainty recognition and prompt-sensitive activation. In the first stage, we guide the LLMs to reject unknown questions. In the second stage, we recover the decreased performance in question-answering by incorporating designed causal instructions. Our goal with this method is to improve the model’s ability to identify and acknowledge areas of uncertainty. The results of our experiment show that the uncertainty-sensitive tuning method greatly enhances the performance of the Llama2-chat-7B Touvron et al. (2023) model. In particular, it brings about a significant 34.7% improvement in addressing questions that involve knowledge gaps compared to the original model. Furthermore, our approach surpasses GPT-4-1106-preview OpenAI (2023), demonstrating a 9.4% increase in overall performance.
In conclusion, our contributions are:
  • We investigate why LLMs tend to ignore essential instructions on this dataset and attribute it to weak sensitivity to uncertainty.
  • We construct a novel dataset for uncertainty-aware assessment, which assisted in evaluating the model’s awareness of knowledge gaps.
  • We propose a novel two-stage fine-tuning paradigm for instructing the model to reject the questions without sufficient context, mitigating the probability of hallucination in this situation.
2 Related Work
In this section, we analyze the former works about hallucinations and instruction datasets for training.
2.1 Uncertainty In Hallucinations
Although the LLM has demonstrated strong performance in downstream tasks by generalizing and leveraging encoded knowledge within the parameters Pu and Demberg (2023); Zhang et al. (2023b), the uncertainty of such knowledge can also mislead models to generate untrustworthy outputs Yu et al. (2023a); Ye et al. (2023); Manakul et al. (2023). Generally, the uncertainty comes from training data and overestimation Zhang et al. (2023a). Research shows that models tend to mimic the output in the training set Kang and Hashimoto (2020), leading to hallucinations that generate reasonable answers for insufficient question-context pairs. Furthermore, models could be overconfident in their capacities and fail to identify unknown questions Yin et al. (2023); Ren et al. (2023); Kadavath et al. (2022).
There are studies focusing on recognizing hallucinations through uncertainty measurement and mitigation. Lu et al. (2023) propose a functional uncertainty measurement approach and conclude that a correlation exists between the uncertainty and the accuracy. CAD Shi et al. (2023) proposes a contrastive method for measuring the uncertainty of generated knowledge, restricting models to be context-awarded by amplifying output probabilities when the context is provided. SelfCheckGPT Manakul et al. (2023) utilizes multiple sampling to identify and exclude uncertain information. Our proposed method instructs the model to reject questions that exceed its existing knowledge boundaries. This helps the model mitigate the hallucinations of uncertainty during the initial training process.
2.2 LLMs With External Knowledge
Existing research has demonstrated that incorporating up-to-date, relevant knowledge in the prompt can effectively reduce fact-conflicting hallucination Vu et al. (2023); Lewis et al. (2020). Zhou et al. (2023) proposes multiple formats for prompt integration of the additional knowledge. Liu et al. (2022) suggests fine-grained hallucination detection through providing token-level ground truths. However, those works are based on the assumption that the provided knowledge is always sufficient for answering. To address the issue, most works choose to employ LLMs to further determine the consistency between the generated answers and contexts. Shen et al. (2023) takes an LLM classifier appended with a generator, while PURR Chen et al. (2023) appends an LLM denoiser, for detecting hallucinations in a generation. Self-RAG Asai et al. (2023) follows a similar approach, utilizing LLMs to screen the provided context, avoiding the disruptions of irrelevant information. However, models struggle to accurately determine whether the provided knowledge is sufficient for answering, especially when the domains of query and context are similar.
Besides, some research has found that providing the LLMs with ’unknown’ knowledge will decrease performance, potentially as it encourages hallucination Lee et al. (2024). Thus, we need to equip the LLM with the ability to know the ’unknown’, which is the main contribution of this work.
2.3 Instruction Dataset for Training
Since Large Language Model (LLM) alignment requires a significant amount of training data, there is a trend toward LLMs synthesizing instruction data to improve their performance. For example, Self-Instruct Wang et al. (2023) initially proposed generating a variety of instructions using ChatGPT. To improve the query’s complexity in different dimensions, WizardLM Xu et al. (2023) uses five prompts, including depth search and with search. On the other hand, AttrPrompt Yu et al. (2023b) does not rely on class-conditional prompts but generates multiple instructions from a feature perspective. However, the majority of these methods concentrate on enhancing the quality of answers by exploring a variety of questions when a definite answer is available, rather than addressing what to do when a question lacks a definitive answer. Thus, R-Tuning Zhang et al. (2024) instructs LLMs to reject unknown questions, but they only focus on rejecting out-of-domain problems. After fine-tuning the rejection questions, we discover that LLMs experience a decrease in answer correctness, necessitating performance improvement. Therefore, in this study, we propose a two-step training method to enable language models (LLMs) to reject the ’unknown’ question while still maintaining their performance on other tasks.
3 Method
In this section, our focus lies on our proposed uncertainty-sensitive tuning. We introduce uncertainty recognition as a two-stage training process. Figure 2 shows the overview and two training steps for the proposed method.
Figure 2: Illustration of the Uncertainty-Sensitive Tuning. We divide the training procedure into two stages. The orange model is not finetuned, the yellow model is finetuned after stage 1, while the green model is our uncertainty-sensitive tuned model. The text in the green background dialog box is a task-oriented instruction used in the fine-tuning process. Uncertainty-Recognition Tuning (Stage 1): teaching the model to know the unknown. Prompt-Sensitive Tuning (Stage 2): teaching the model to follow the instructions.
3.1 Question Definition
Our research problem setting is similar to the open-book question-answering task. This task aims to generate an answer $a$ based on three inputs: $i_{task}$, $q$, and $c$. Here, $i_{task}$ represents the designed task instructions, $q$ is the specific question, and $c$ is the given context. This process can be expressed as:
$a=G(i_{task},q,c)$
Where $G$ is the model that takes the instructions ($i_{task}$), the query ($q$), and the context ($c$) as inputs and generates the answer $a$ as output.
In $i_{task}$, we will explicitly instruct the model to not utilize knowledge outside the provided context using the prompt "Your answer must not use any additional knowledge that is not mentioned in the given contexts". This process relies on the generative model $G$ to evaluate whether the context $c$ is sufficient to answer the question $q$. Based on this assessment, $G$ either generates an appropriate response ($a$) or acknowledges the insufficiency of $c$.
3.2 Preliminary Exploration On The Unknown Question
Given the research question, "How to instruct the LLM to know the unknown questions?", we preliminary finetune the Llama2-Chat-7B Touvron et al. (2023) on the questions that exceed the provided context (detailed experiment settings are in Section 4) and use the vanilla Llama2-Chat-7B as the baseline. Table 1 illustrates the result of this experiment. The $Acc_{known}$ represents the performance of different models on questions with specific answers, and the $Acc_{unknown}$ represents the questions without related context.
Experiments show that LLMs are suffering from two aspects of weakness: First, the vanilla model is unable to accurately reject unknown questions as it is mostly trained with specific answers. Second, the QA performance decreased after training on the unknown question, and even further training on the hotpotQA dataset Yang et al. (2018) failed to mitigate the challenge posed by the unknown question task. Therefore, it is necessary for models to learn how to reject questions they are unfamiliar with without compromising QA performance. To maintain the performance of the question-answering (QA) task while effectively handling unknown questions, we propose a two-step training strategy. The first step involves training the large language model (LLM) to identify and reject questions that fall outside its knowledge domain. This enables the model to recognize when it does not have sufficient information to provide an accurate answer. The second step consists of prompt-sensitive instruction training and review, recovering the decreased performance. This step goes beyond superficial question-answer instruction fine-tuning and comprehensively enhances the model’s ability to understand the intent behind the prompts, even for ’unknown’ questions.
Model $Acc_{known}$ $Acc_{unknown}$ $F1$
Baseline 79.3 58.3 67.2
Unknown Question Finetuned 52.4 84.4 64.6
Unknown Question Finetuned + HotpotQA 77.0 20.9 32.8
Table 1: Results (in %) for the preliminary experiment. The baseline is the vanilla Llama2-Chat-7B. We report the accuracy of different questions in the evaluation dataset (detailed in section 4).
3.3 Stage 1: Uncertainty-Recognition Tuning
The first stage is uncertainty-recognition tuning, which focuses on tuning a model to have an accurate awareness of its lack of knowledge. Shown in figure 2, we form this task as a binary classification problem. Questions are classified into two types: known questions and unknown questions. A known question is characterized by sufficient contextual information to provide an answer. In contrast, an unknown question is characterized by insufficient contextual information to provide a response. Models are required to detect whether the provided context is sufficient to answer the question. Specifically, given a dataset $D=\{(q_{1},c_{1}),(q_{1},{c^{\prime}}_{1})...(q_{n},c_{n}),(q_{n},{c^{\prime}}_{n}))\}$ consisting of $n$ positive question-context pairs and $n$ negative ones, we finetune the LLM to do the classification; only "Sufficient" or "Insufficient" is accepted as an answer. The instruction for uncertainty-recognition tuning is provided in Appx. A.2.
3.4 Stage 2: Prompt-Sensitive Tuning
The model obtained through uncertainty-recognition tuning can recognize the boundary of the given context. However, table 1 shows that the model may be oversensitive to unknown questions. After stage 1, the model indeed learned to answer the unknown questions accurately. However, the performance of questions with specific answers decreased by nearly 27%. We hypothesize that models fine-tuned on unknown questions suffer from a prompt-sensitivity reduction problem and may break some instructions that would ideally lead to better generation. We categorize the factors influencing the response in the instructions into causal and non-causal factors. Causal factors are important as they directly affect the response, whereas non-causal factors have minimal impact on the result. The control conditions in the instruction, such as the word number of responses, are a typical causal factor. The extra instruction, such as ’give me a reference if inputs have a file name’, is a non-causal factor because it may not change the answer. During the training process, LLM may ignore this type of instruction because it doesn’t always directly contribute to the answers. But the non-causal factor is equally important to the question. For example, the introduction of a book might seem unrelated to the main content, leading readers to underestimate its value. However, it plays a crucial role in enhancing their understanding of the content.
Based on this, we further propose prompt-sensitive tuning, aiming at guiding the model to fulfill all the instructions in the prompt. As shown in Fig. 2, prompt-sensitive tuning consists of two sub-tasks: additional causal instructions and instruction review. We use GPT-4 to synthesize the required data and distill it into the target model.
Causal Instruction Synthesis: By instructing GPT-4 to generate controlling requirements guaranteed to affect the final response, we obtain additional causal instructions. Then, to improve the model’s instruction sensitivity, we randomly insert one causal instruction into the original QA instruction. For example, limit the number of words, tense, or format of the output. The prompt is presented in Appx. A.3. Then, we prompt GPT-4 to respond with new instructions and record the conversation data to fine-tune the smaller model.
Review Instruction Synthesis: The instruction review module aims at using the model itself to verify that all instructions are completed. The model will recursively regenerate until it gets a perfect answer by utilizing customized prompts recorded in Appx. A.4. The process of the instruction review is illustrated in Algo. 1.
Data: context $c$, query $q$, instructions $i$
1
2fulfilled = False;
3 while not fulfilled do
4       answer, checking = generate($c$, $q$, $i$);
5       if "<not fulfilled>" not in checking then
6            fulfilled = True;
7       end if
8
9 end while
10
Result: answer
Algorithm 1 Instruction Review Module
As shown in Fig. 2, given a question answering dataset $\{(q_{1},c_{1}),(q_{2},c_{2})...(q_{n},c_{n})\}$ and additional causal instructions, the entire process could be formulated as $a=R(G(i_{task}+i_{casual},q,c))$, where $i_{casual}$ is a randomly chosen casual instruction and $i_{task}$ is the original task description. $R$ is the loop function for instruction reviewing. Meanwhile, we also prompt GPT-4 and record the conversation data from the loop to fine-tune the smaller model.
4 Experiments
In this section, we explain the construction of our instruction finetuning dataset and the corresponding experiments. We evaluate the proposed method among mainstream LLMs and present the results in the subsequent subsections.
QA Uncertainty-Recognition
Postfix
Uncertainty-Recognition
Model $Acc_{known}$ $Acc_{unknown}$ $F1$ $Acc_{known}$ $Acc_{unknown}$ $F1$
GPT-4 Turbo 79.6 83.6 81.6 88.9 78.3 83.3
GPT-3.5 Turbo 82.1 51.8 63.5 97.0 33.4 49.7
Vicuna-7B v1.5 74.6 43.8 55.2 93.5 14.3 24.8
Self-RAG-7B 67.9 48.1 56.3 46.0 74.9 57.0
Llama2-Chat-7B (base) 79.3 58.3 67.2 85.2 29.5 43.9
Uncertainty-Recognition Tuned 52.4 84.4 64.6 81.3 84.0 82.6
Uncertainty & HotpotQA Tuned 77.0 20.9 32.8 87.1 66.7 75.5
Uncertainty-Sensitivity Tuned 79.7 93.0 85.8 88.0 66.0 75.4
Table 2: Results (in %) for mainstream and our finetuned LLMs on QA uncertainty-recognition assessment and postfix uncertainty-recognition assessment. The overall best results are highlighted in bold. The results obviously higher and lower than the baseline are highlighted in green and orange, respectively.
4.1 Data Construction
Following the proposed two-stage framework, we synthesize two new instruction datasets for LLM finetune. Besides, we construct a prefix uncertainty-recognition evaluation dataset to better evaluate the LLMs’ recognition of not-knowing.
Uncertainty-Sensitive Tuning Dataset In uncertainty-sensitive tuning, there are two distinct types of instruction datasets: the binary uncertainty-recognition tuning dataset and the prompt-sensitive fine-tuning dataset. For the binary uncertainty-recognition tuning dataset, we sample 646 examples from the ASQA dataset Stelmakh et al. (2022), which consists of ambiguous question-context pairs. Each pair is divided into multiple sub-questions with their corresponding contexts. For example, one pair may discuss the discovery of the photoelectric effect in 1887, while another pair may cover the theoretical development of the photoelectric effect in 1905. To generate the unknown questions, we shuffle these pairs, reassigning the questions to different but related contexts. The ambiguity between question-context pairs introduces minor deviations that challenge the model’s ability to handle uncertainty, differentiating our method from other works. The prompt-sensitive fine-tuning dataset is based on HotpotQA Yang et al. (2018), a dataset designed for multi-hop reasoning question answering. To adapt this dataset for fine-tuning with additional causal instructions, we generate and manually select the most effective 28 causal instructions. Then, we integrate these instructions with 300 randomly selected samples from HotpotQA. Finally, all the generated datasets are used for fine-tuning. Here is a demonstration of the additional prompt in uncertainty-recognition tuning:
  • Uncertainty-Recognition Tuning: You must only answer either ’Sufficient’ or ’Insufficient’ without any other output
Uncertainty-Recognition Benchmark To comprehensively evaluate the model’s cognitive ability to recognize knowledge gaps, we construct a testbed using the ASQA Stelmakh et al. (2022)dataset, ensuring no overlap with the training data. During the evaluation, two types of designed evaluation templates are available: the QA uncertainty recognition template and the postfix uncertainty recognition template. For the QA uncertainty recognition, we instruct the model to either answer the question or reject it with ’Not Provided’. We generate 3320 test data for this type of template. For postfix uncertainty recognition, we instruct the LLMs to answer the question before classifying it. We include 646 test samples for this template. Here is a demonstration of the used prompt templates:
  • QA Uncertainty-Recognition: If the context is not sufficient to answer the question, please answer it with ’Not Provided’
  • Postfix Uncertainty-Recognition: You must append either ’<Sufficient>’ or ’<Insufficient>’ after your answer.
4.2 Experiment Setting
Training Details. Our experiments are conducted on mainstream LLMs, including Llama2-Chat-7B Touvron et al. (2023), GPT-4 Turbo OpenAI (2023) , GPT-3.5 Turbo OpenAI (2023), Vicuna-7B v1.5 Zheng et al. (2024a) and Self-RAG-7B Asai et al. (2023). Taking Llama2 as the baseline, we further finetune it on a single GTX3090 GPU by LLaMA-Factory Zheng et al. (2024b). Training with Lora Hu et al. (2021) in a rank of 8, with a batch size of 4 and a learning rate of 5e-5. This research integrates the instruction-based prompt and attributed prompt which is proved to be effective for mitigating hallucination Zhou et al. (2023), as provided in Appx. A. We fine-tuned three different versions of the model, each using distinct instruction datasets. The uncertainty-recognition-tuned model was fine-tuned using the proposed uncertainty-recognition-tuning dataset. The Uncertainty & HotpotQA tuned model was fine-tuned with both the proposed uncertainty-recognition tuning dataset and the HotpotQA dataset. Lastly, the uncertainty-sensitivity tuned model utilized a combination of the uncertainty-recognition tuning dataset and the prompt-sensitivity tuning dataset.
Evaluation Metric. We use $Acc_{known}$ to demonstrate the accuracy of questions with specific answers and $Acc_{unknown}$ represents the accuracy of unknown questions.
4.3 Results
Table 2 presents the results from our evaluation of mainstream large language models (LLMs). The evaluation of question-answering (QA) uncertainty reveals that most large language models (LLMs) struggle to accurately identify unknown questions, with only around 50% accuracy. However, GPT-4 stands out as a notable exception, outperforming the second-best model, Llama2, by a significant margin of 25.3%. Llama2 ranks highest among the remaining models, even surpassing GPT-3.5, despite the latter having more parameters. But there is still a substantial performance gap between the GPT-4 and the other models. The ongoing experiments aim to investigate the reasons behind this disparity. The postfix uncertainty-recognition assessment results, shown in the right column of Table 2, indicate a strong positive correlation between the two evaluation templates. The analysis also highlights that different models respond differently to insufficient information. Models fine-tuned on dialogue tasks tend to overly rely on and trust the given information. Self-RAG, which is fine-tuned for question-answering tasks with unknown questions, demonstrates a strong ability to reject questions it cannot answer. However, despite Self-RAG’s impressive performance in the uncertainty-recognition assessment, it remains difficult to train the model to effectively acknowledge unknown information.
Our fine-tuned model, which has been tuned for uncertainty recognition, performs comparably to GPT-4 in tests of uncertainty recognition and shows a 25.9% improvement in responding to context compared to the baseline model. However, despite outperforming GPT-4 in terms of unknown questions ($Acc_{unknown}$), this increased awareness of uncertainty leads to a decrease in the model’s original question-answering capabilities. Further fine-tuning on the HotpotQA dataset resulted in the model being less reliable in following instructions and sometimes disregarding its knowledge limitations. Among the models tested, the uncertainty-sensitivity tuned model ranks the highest, achieving an F1 score of 85.8, which is 4.2 points higher than GPT-4 and 18.6 points above the baseline (as shown in Figure 2). This impressive performance is due to the model’s ideal balance between answering questions and acknowledging when it lacks the necessary information to provide an answer. It achieves a remarkable 93.0% accuracy on unknown questions, the highest among all mainstream LLMs, while maintaining a 79.7% accuracy on known questions. Moreover, the model effectively integrates new information without compromising its existing knowledge base, perfectly retaining the knowledge it acquired from cognition fine-tuning.
4.4 Ablation Study
4.4.1 Experiment Setting
In order to further quantify the effects of uncertainty-recognition tuning, HotpotQA tuning and prompt-sensitive tuning, three additional variants have been built for ablation experiments.
  • The HotpotQA finetuning utilizes 300 samples with traditional question answering prompts provided in Appx. A.1.
  • The prompt-sensitive tuning bases on the 300 samples mentioned in Section 3.4.
  • The Uncertainty & HotpotQA tuned model further finetunes the uncertainty-recognition tuned model on HotpotQA dataset.
4.4.2 Result Analysis
Table 2 demonstrates that the reasons for LLMs’ poor performance in rejecting the unknown questions come from two aspects: weak uncertainty-recognition capacity and the prompt-sensitivity reduction problem.
Weak Uncertainty-Recognition Capacity: Our ablation study, as shown in Fig. 3, illustrates that enhancing the model’s ability to recognize uncertainty significantly boosts performance on unknown questions. Specifically, the uncertainty-recognition tuning resulted in a marked 16.9% increase in ${Acc}_{unknown}$. The models that have not been trained on unknown tasks exhibit poor performance. We call this weak uncertainty recognition. The problem of weak uncertainty-recognition capacity in LLMs comes from their fundamental training process to generate specific responses. In scenarios where LLMs encounter questions outside their training distribution or scope of knowledge, they often fail to adequately recognize uncertainty. This leads to inappropriate or misleading answers.
Prompt-sensitivity reduction problem: The reduction in known questions is another critical issue highlighted by the ablation study. As models are further tuned to handle specific tasks like QA finetuning or uncertainty recognition, they can become less responsive to variations in the input prompt, a phenomenon we term "prompt-sensitivity reduction." This reduced sensitivity can impair the model’s ability to adapt to the slight but significant differences in how questions are phrased or presented, leading to what we describe as cognitive inaccessibility. From Fig. 3, it is evident that although simple QA finetuning attempts to mitigate the negative effects of over-tuning for uncertainty recognition, it does not fully address the prompt-sensitivity issue. The model continues to struggle with lower overall accuracy because it cannot effectively adapt to the diverse ways in which users might phrase similar questions. Thus, the model fails to fully leverage the context of the prompt, resulting in inaccurate or relevant answers. With our prompt-sensitive tuning, we can successfully recover the performance of other questions. Even if we do not fine-tune the model for the unknown task, our prompt-sensitive tuning can boost the performance of the task with specific answers, which is the highest accuracy in the table.
QA Uncertainty-Recognition
Model $Acc_{known}$ $Acc_{unknown}$ $F1$
Llama2-Chat-7B (baseline) 79.3 58.3 67.2
Uncertainty-Recognition Tuned 52.4 84.4 64.6
HotpotQA Tuned 77.5 58.3 66.5
Prompt-Sensitive Tuned 84.8 59.0 69.6
Uncertainty & HotpotQA 77.0 20.9 32.8
Uncertainty-Sensitive Tuned 79.7 93.0 85.8
Table 3: Results of ablation. The overall best results are highlighted in bold. The results obviously higher and lower than the baseline are highlighted in green and orange, respectively.
Figure 3: Comparison of different models with different training dataset. A higher score means better performance.
5 Conclusion
In this paper, we explore a common issue in language model applications where insufficient context in prompts leads to the generation of plausible yet incorrect responses, a phenomenon known as hallucinations. Firstly, our research identifies that large language models (LLMs) often fail to recognize when they lack adequate information to accurately respond to a query. We primarily attribute this problem to the fact that the models lack the ability to recognize when the provided context is insufficient for the posed queries. Second, during the training process with questions without a specific answer, the performance of questions with answers decreased. To counter these issues, we introduce a novel training methodology termed "uncertainty-sensitive tuning."  This is a two-stage training framework. In the first stage, we guide the LLM to reject unknown questions. In the second stage, we recover the decreased performance in QA tasks by incorporating designed causal instructions. This approach significantly enhances the Llama2-chat-7B model’s ability to handle queries more reliably, reducing the occurrence of hallucinations. Our approach differs from previous works in that it focuses more directly on refining the model’s response behavior through targeted fine-tuning, thereby enhancing its performance in practical scenarios. By open-source this framework and model, we believe that it will highlight a new direction for the automatic synthesis instruction data set, which is not only focused on the diversity of the data but also the reduction of hallucinations during training.
Limitations
In this work, we have two promising directions to refine in the future. The first limitation is that when tuned with the new datasets, which contain a limited number of unknown questions, the LLM experiences a long-tail problem. Therefore, to effectively tailor our proposed two-stage, uncertainty-sensitive method to specific datasets, it requires further tuning. Second, we face a limitation in that we did not probe or analyze the parametric knowledge learned by Llama2 during its initial pre-training phase. Our fine-tuning dataset may overlap with Llama2’s pre-training data. To address these challenges, our future work aims to explore existing methods that measure model uncertainty by monitoring internal parameters Lu et al. (2023). By quantifying the model’s uncertainty on different inputs, we can gain insights into its knowledge gaps and long-tail weaknesses. This knowledge could then lead to more targeted fine-tuning strategies or engineering methods that make the fine-tuned LLM work better on different types of queries.
Ethics Statement
Our benchmark and datasets are constructed from public datasets, and uncertainty-sensitive tuning dataset further introduces GPT-4 to refine the dataset, thus inherent biases may exist. However, the techniques utilized in this work do not introduce additional biases beyond what the data has presented.
Appendix A Instructions
A.1 Question Answering
Question Answering Task: You need to do the Question Answering for the following query. I will give a question and several contexts. Based on the given contexts, give an answer to the question. Your answer must not using any additional knowledge that is not mentioned in the contexts. If the context is not sufficient to answer the question, please answer it with ’Not Provided’ QUERY: $q$ CONTEXT: $c$ ANSWER:
A.2 Uncertainty-Recognition
Cognition Assessment Task: You need to do the Cognition Assessment for the following query. I will give a query and a related context about the query. Your task is to judge whether the context is sufficient to answer the query. Assessment: You must append either ’<Sufficient>’ or ’<Insufficient>’ after your answer. Finetuning: You must only answer either ’Sufficient’ or ’Insufficient’ without any other output. Here is the example. QUERY: What happened to Jay when he got old? CONTEXT: Jay Chou was the most famous singer in China when he was young, releasing many nostalgic albums and songs that are memorable to middle-aged people today. ANSWER: Assessment: Jay Chou was the most famous singer in China.<Insufficient> Finetuning: Insufficient Here is the provided information that you need to accomplish follow the provided example: QUERY: $q$ CONTEXT: $c$ ANSWER:
A.3 Additional Casual Instruction Generation
Your task is to provide various instructions for the questions answering task. The questions answering task provides a context and a query. e.g. "Context: XXX Query: XXX Answer:". And your task is to add some specific requirement to the answer. e.g. "The answer must be all in upper case", "There should be no punctuation in the answer". The added instruction should be general to the query. You should generate hundreds of such instructions.
A.4 Prompt-Sensitive Finetuning
You should check whether your answer aligned the requirement by generating a Checking part, checking each sentence of the above instruction, with either <fulfilled> or <not fulfilled> mark behind the sentence, indicating whether the requirement is fulfilled or not. If there is <not fulfilled> mark behind the sentence, you must modify your answer again to fulfill the requirement, by appending a new ANSWER and CHECKING part. Here is an example for this task: e.g. Question Answering Task Requirements: You need to do the Task Prompt for the following query and context. Ensure the response is written in the past tense. QUESTION: Who is Jack Chen? CONTEXTS: People saying that Jack Chen is a famous singer in China. ANSWER: Jack Chen was a famous singer in China. CHECKING: Question Answering Task: You need to do the Task Prompt for the following query and context.<fulfilled>Ensure the response is written in the past tense.<not fulfilled> ANSWER: Jack Chen was a famous singer in China. CHECKING: Question Answering Task: You need to do the Task Prompt for the following query and context.<fulfilled>Ensure the response is written in the past tense.<fulfilled> Here is the information of your task: {Question Answering Instruction}
A.5 Trustworthy Question Answering for Benchmark
Trustworthy Question Answering Task: You need to utilize the ability learnt during both the Question Answering Task and Cognition Assessment Task. And only provide the answers which are sufficiently supported by the context, otherwise provide ’Not Provided’ I will give a question and several context texts about the question. Based on the given contexts, give an answer to the question. Your answer must not using any additional knowledge that is not mentioned in the given contexts. If the context is not sufficient to answer the question, please answer it with ’Not Provided’ QUERY: $q$ CONTEXT: $c$ ANSWER:
Appendix B Example of Proposed Benchmark
Question: Who discovered and developed an explanation for the
photoelectric effect in 1887?
Positive Context: This is a passage about Photoelectric effect:
Light, and especially ultra-violet light, discharges negatively electrified
bodies with the production of rays of the same nature as cathode rays.
Under certain circumstances it can directly ionize gases. The first of
these phenomena was discovered by Heinrich Hertz and Wilhelm
Hallwachs in 1887. The second was announced first by Philipp Lenard in 1900.
Negative Context: This is a passage about Photoelectric effect:
In 1905, Einstein proposed an explanation of the photoelectric effect using a
concept first put forward by Max Planck that light waves consist of tiny
bundles or packets of energy known as photons or quanta.