### v2

# Know the Unknown: An Uncertainty-Sensitive Method for LLM Instruction Tuning

Abstract
Large language models (LLMs) have demonstrated remarkable capabilities but still face challenges such as hallucinations. One potential reason for hallucinations is the lack of relevant knowledge or context. Thus, a promising solution involves instructing LLMs to respond with "I do not know" when a question falls outside their knowledge domain or the provided context. However, in this work, we observed that LLMs struggle to admit their lack of knowledge, primarily due to existing instruction datasets designed to encourage specific answers. To improve models’ capability to recognize the boundaries of their knowledge, we propose a novel approach called uncertainty-sensitive tuning. This method involves two-stage training designed for uncertainty recognition and prompt-sensitive activation. In the first stage, we guide the LLM to reject unknown questions. In the second stage, we force the model to follow the instructions by incorporating designed causal instructions. The experimental results demonstrate that our proposed uncertainty-sensitive tuning method enhance the model’s ability to identify areas of uncertainty. Specifically, it achieves a substantial improvement of up to 34.7% in handling questions involving knowledge gaps compared to the original model. Moreover, our finetuned models even outperform GPT-4, exhibiting an overall performance improvement of up to 4.2%.  [^1] The model and code are avaliable at https://github.com/JiaqiLi404/TrustworthyRAG.
Know the Unknown: An Uncertainty-Sensitive Method for LLM Instruction Tuning
Jiaqi Li1, Yixuan Tang1, Yi Yang1 1The Hong Kong University of Science and Technology Correspondence: imyiyang@ust.hk
1 Introduction
The advancement of large language models (LLMs) has shown great potential for various tasks Brown et al. (2020); Wei et al. (2022); Joshi et al. (2017). However, there are still some issues, such as hallucinations Huang et al. (2023); Ji et al. (2023); Ye et al. (2023). Hallucination is a phenomenon in which LLMs generate a response that is either factually incorrect or nonsensical Maynez et al. (2020). Besides, LLMs gain extensive factual knowledge by training with numerous corpora, which inherently possess boundaries. When faced with unfamiliar information, the LLM’s response tends to be nonsense or outdated Huang et al. (2023).
Figure 1: The response of different models facing two types of questions. Known Question: The question has specific answers. Unknown Question: The question falls outside the provided context.
To bridge the knowledge gap, there are two main approaches. The first approach is to add new knowledge to the model, such as further fine-tuning Liu et al. (2023b); Gao et al. (2023); Liu et al. (2023a). The second approach is to equip the LLMs with an external database and use retrieval-augmented generation Es et al. (2023). However, the existing instruction dataset focuses on teaching LLMs to answer questions with a specific answer, making models incapable of responding "I don’t know" Zhang et al. (2023a), which will also increase the tendency for hallucination. To address this issue, many researchers Zhang et al. (2024); Cole et al. (2023); Yu et al. (2024) suggest prompting LLMs to admit their lack of knowledge. However, by doing this, LLMs may ignore important instructions contained in the prompts Kasai et al. (2023); Li et al. (2023); Si et al. (2023). As shown in Fig. 1, LLMs may still utilize knowledge outside the provided corpus even with explicit instructions, such as "If the context is not sufficient to answer the question, please answer it with ’Not Provided’.".
This paper is motivated by the findings that when there are tiny but significant discrepancies between the knowledge and the question, LLMs tend to be hallucinating Huang et al. (2023); Shuster et al. (2021). Thus, we focus on the situation in which the provided context is insufficient. Initially, we explore why LLMs struggle to acknowledge their limitations in knowledge. We hypothesize that this issue derives from the models’ perceptual capabilities and introduces uncertainty-aware assessment. We utilize a dataset featuring known and unknown questions, where the unknown questions are crafted with insufficient context for answering questions. The dataset is generated using the special ambiguous query-context pairs in the ASQA Stelmakh et al. (2022). We pair each question with a slightly insufficient context, such as describing different dates or objects while keeping the rest of the content reasonable. Compared to the other works Li et al. (2022); Chen et al. (2023), the slight difference could extremely confuse and guide the LLMs to hallucinate. This dataset, called the uncertainty-recognition benchmark, serves as the foundation for our performance assessments.
To enhance the capabilities of LLMs in recognizing the limits of their knowledge, we introduce a novel training framework called uncertainty-sensitive tuning. This approach contains a two-stage training process specifically tailored for uncertainty recognition and prompt-sensitive activation. In the first stage, we guide the LLMs to reject unknown questions. In the second stage, we recover the decreased performance in question-answering by incorporating designed causal instructions. Our goal with this method is to improve the model’s ability to identify and acknowledge areas of uncertainty. The two-stage strategy is necessary for addressing the overall performance degradation when fine-tuning on normal question answering (QA) datasets, which are filled with known queries. We attribute such degradation to the model’s weak sensitivity to invariant prompts, which is improved in the second stage. The results of our experiment show that the uncertainty-sensitive tuning method greatly enhances the performance of the mainstream LLMs. In particular, it brings about a significant 34.7% improvement in addressing questions that involve knowledge gaps compared to the original Llama2-7B-Chat Touvron et al. (2023). Furthermore, our approach surpasses GPT-4-1106-preview OpenAI (2023), demonstrating an overall performance increase of up to 4.2%.
In conclusion, our contributions are:
  • We investigate why LLMs could break the limitations form essential instructions after traditional QA tuning, and attribute it to the weak sensitivity to invariant prompts.
  • We construct a novel dataset for uncertainty-recognition assessment, which assisted in evaluating the model’s awareness of knowledge gaps.
  • We propose a novel two-stage fine-tuning paradigm for instructing the model to reject the questions without sufficient context, mitigating the probability of hallucination in this situation.
2 Related Work
In this section, we analyze the former works about hallucinations and instruction datasets for training.
2.1 Uncertainty in Hallucinations
Although the large language models (LLMs) has demonstrated strong performance in downstream tasks by generalizing and leveraging encoded knowledge within the parameters Pu and Demberg (2023); Zhang et al. (2023b), the uncertainty of such knowledge can also mislead models to generate untrustworthy outputs Yu et al. (2023a); Ye et al. (2023); Manakul et al. (2023). Generally, the uncertainty comes from training data and overestimation Zhang et al. (2023a). Research shows that models tend to mimic the output in the training set Kang and Hashimoto (2020), leading to hallucinations that generate reasonable answers for insufficient question-context pairs. Furthermore, models could be overconfident in their capacities and fail to identify unknown questions Yin et al. (2023); Ren et al. (2023); Kadavath et al. (2022).
There are studies focusing on recognizing hallucinations through uncertainty measurement and mitigation. Lu et al. (2023) propose a functional uncertainty measurement approach and conclude that a correlation exists between the uncertainty and the accuracy. CAD Shi et al. (2023) proposes a contrastive method for measuring the uncertainty of generated knowledge, restricting models to be context-awarded by amplifying output probabilities when the context is provided. SelfCheckGPT Manakul et al. (2023) utilizes multiple sampling to identify and exclude uncertain information. Our proposed method instructs the model to reject questions that exceed its existing knowledge boundaries. This helps the model mitigate the hallucinations of uncertainty during the initial training process.
2.2 LLMs with External Knowledge
Existing research has demonstrated that incorporating up-to-date, relevant knowledge in the prompt can effectively reduce fact-conflicting hallucination Vu et al. (2023); Lewis et al. (2020). Zhou et al. (2023) proposes multiple formats for prompt integration of the additional knowledge. Liu et al. (2022) suggests fine-grained hallucination detection through providing token-level ground truths. However, those works are based on the assumption that the provided knowledge is always sufficient for answering. To address the issue, most works choose to employ LLMs to further determine the consistency between the generated answers and contexts. Shen et al. (2023) takes an LLM classifier appended with a generator, while PURR Chen et al. (2023) appends an LLM denoiser, for detecting hallucinations in a generation. Self-RAG Asai et al. (2023) follows a similar approach, utilizing LLMs to screen the provided context, avoiding the disruptions of irrelevant information. However, models struggle to accurately determine whether the provided knowledge is sufficient for answering, especially when the domains of query and context are similar.
Besides, some research has found that providing the LLMs with ’unknown’ knowledge will decrease performance, potentially as it encourages hallucination Lee et al. (2024). Thus, we need to equip the LLM with the ability to know the ’unknown’, which is the main contribution of this work.
2.3 Instruction Dataset for Training
Since LLM alignment requires a significant amount of training data, there is a trend toward LLMs synthesizing instruction data to improve their performance. For example, Self-Instruct Wang et al. (2023) initially proposed generating a variety of instructions using ChatGPT. To improve the query’s complexity in different dimensions, WizardLM Xu et al. (2023) uses five prompts, including depth search and with search. On the other hand, AttrPrompt Yu et al. (2023b) does not rely on class-conditional prompts but generates multiple instructions from a feature perspective. However, the majority of these methods concentrate on enhancing the quality of answers by exploring a variety of questions when a definite answer is available, rather than addressing what to do when a question lacks a definitive answer. Thus, R-Tuning Zhang et al. (2024) and Cole et al. (2023) instructs LLMs to reject unknown questions, but they only focus on rejecting out-of-domain problems. After fine-tuning the rejection questions, we discover that LLMs experience a decrease in answer correctness, necessitating performance improvement. Therefore, in this study, we propose a two-step training method to enable language models (LLMs) to reject the ’unknown’ question while still maintaining their performance on other tasks.
3 Uncertainty-Sensitive Tuning
In this section, our focus lies on our proposed uncertainty-sensitive tuning. Figure 2 shows the overview and two stages for the proposed method.
Figure 2: Illustration of the Uncertainty-Sensitive Tuning. We divide the training procedure into two stages. The orange model is not fine-tuned, the yellow model is fine-tuned after stage 1, while the green model is our uncertainty-sensitive tuned model. The text in the green background dialog box is a task-oriented instruction used in the fine-tuning process. Uncertainty-Recognition Tuning (Stage 1): teaching the model to know the unknown. Prompt-Sensitive Tuning (Stage 2): teaching the model to follow the instructions.
3.1 Question Definition
Our research problem setting is similar to the open-book question-answering (QA) tasks. This task aims to generate an answer $a$ based on three inputs: $i_{task}$, $q$, and $c$. Here, $i_{task}$ represents the designed task instructions, $q$ is the specific question, and $c$ is the given context. The generation process $G$ can be expressed as:
$a=G(i_{task},q,c)$
In $i_{task}$, we will explicitly instruct the model to not utilize knowledge outside the provided context using the prompt "Your answer must not use any additional knowledge that is not mentioned in the given contexts". This process relies on the model $G$ to evaluate whether the context $c$ is sufficient to answer the question $q$. Based on this assessment, $G$ either generates an appropriate response ($a$) or acknowledges the insufficiency of $c$.
3.2 Preliminary Exploration on the Unknown Question
Given the research question, "How to instruct the large language models (LLMs) to know the unknown questions?", we construct a valuable and rare dataset that balances known and unknown questions as benchmark, and preliminary fine-tune the Llama2-7B-Chat Touvron et al. (2023) on the dataset with ground truth hidden (detailed experimental settings are in section 4). Table 1 illustrates the result of this experiment. The $Acc_{known}$ represents the performance of normal question answering task, and the $Acc_{unknown}$ evaluates the capacity of admitting the unknown.
Experiments show that LLMs are suffering from two aspects of weakness: First, the vanilla model is unable to accurately reject unknown questions as it is mostly trained with specific answers. Second, the QA performance decreased after training on the unknown question, and even further training on the hotpotQA dataset Yang et al. (2018) failed to mitigate the challenge posed by the unknown question task. Therefore, it is necessary for models to learn how to reject unknown questions without compromising QA performance. Hence, we propose a two-step training strategy. The first step involves training the large language model to identify and reject questions that fall outside its knowledge domain. This enables the model to recognize when it does not have sufficient information to provide an accurate answer. The second step proposes using prompt-sensitive training and review instead of traditional QA tuning, recovering the decreased performance. This step goes beyond superficial QA tuning and comprehensively enhances the model’s ability to understand the intent behind the prompts.
Model $Acc_{known}$ $Acc_{unknown}$ $F1$
Baseline 79.3 58.3 67.2
Balance Dataset Finetuned 52.4 84.4 64.6
Balance Dataset Finetuned + HotpotQA 77.0 20.9 32.8
Table 1: Results (in %) for the preliminary experiment. The baseline is the vanilla Llama2-7B-Chat. We report the accuracy of different questions in the benchmark (detailed in section 4).
3.3 Stage 1: Uncertainty-Recognition Tuning
The first stage is uncertainty-recognition tuning, which focuses on tuning a model to have an accurate awareness of its lack of knowledge. For protecting the ground truth in benchmark, we form this task as a binary classification problem as shown in figure 2. Questions are classified into known questions and unknown questions. A known question is characterized by sufficient contextual information to provide an answer. In contrast, an unknown question is characterized by insufficient contextual information with tiny differences compared to query. Models are required to detect the sufficiency of the provided context. Specifically, given a dataset $D=\{(q_{1},c_{1}),(q_{1},{c^{\prime}}_{1})...(q_{n},c_{n}),(q_{n},{c^{\prime}}_{n}))\}$ consisting of $n$ positive question-context pairs and $n$ negative ones, we fine-tune the LLM to do the classification; only "Sufficient" or "Insufficient" is accepted as an answer. The instruction for tuning is recorded in Appx. A.2.
3.4 Stage 2: Prompt-Sensitive Tuning
The uncertainty-recognition tuned model can recognize the boundary of the given context and answer the unknown questions accurately. However, table 1 shows that the model may be oversensitive to unknown questions. As a result, the performance of questions with specific answers decreased by nearly 27%. In this situation, normal QA tuning would even harm the model’s capacity of rejecting the unknown questions. We hypothesize that models fine-tuned on pure known questions suffer from a prompt-sensitivity reduction problem and may break the instructions about rejecting answering. We categorize the factors influencing the response in the instructions into causal and non-causal factors. Causal factors are important as they directly affect the response, whereas non-causal factors have minimal impact on the result. The control conditions in the instruction, such as the format of responses, are a typical causal factor. The extra instruction, such as "answering with ’Not Provided’ if the context is insufficient", is a non-causal factor when tuning on the known questions, because it may not directly contribute to the answer. During the training process, LLMs may ignore this type of instruction, but the non-causal factor is equally important to the task. For example, the introduction of a book might seem unrelated to the main content, leading readers to underestimate its value. However, it plays a crucial role in enhancing their understanding of the content.
Based on this, we further propose prompt-sensitive tuning, aiming at guiding the model to fulfill all the instructions in the prompt. As shown in Fig. 2, prompt-sensitive tuning consists of two sub-tasks: additional causal instructions and instruction review. We use GPT-4 to synthesize the required data and distill it into the target model.
Causal Instruction Synthesis: By instructing GPT-4 to generate controlling conditions guaranteed to affect the answers, we obtain additional causal instructions. Then, to improve the model’s instruction sensitivity, we randomly insert one causal instruction into the original QA instruction. For example, limit the tense, length, or format of the output. The prompt is presented in Appx. A.3.
Review Instruction Synthesis: The instruction review module employs the model itself to verify the fulfillment of all instructions. The model will recursively regenerate until it gets a perfect answer by utilizing the prompts in Appx. A.4. The process of the instruction review is illustrated in Algo. 1.
Data: context $c$, query $q$, instructions $i$
1
2fulfilled = False;
3 while not fulfilled do
4       answer, checking = generate($c$, $q$, $i$);
5       if "<not fulfilled>" not in checking then
6            fulfilled = True;
7       end if
8
9 end while
10
Result: answer
Algorithm 1 Instruction Review Module
As shown in Fig. 2, given a question answering dataset $\{(q_{1},c_{1}),(q_{2},c_{2})...(q_{n},c_{n})\}$ and additional causal instructions, the entire process could be formulated as $a=R(G(i_{task}+i_{casual},q,c))$, where $i_{casual}$ is a randomly chosen casual instruction and $i_{task}$ is the original task description. $R$ is the loop function for instruction reviewing. We employ GPT-4 and record the conversation data from the loop to fine-tune the smaller model.
4 Experiments
In this section, we explain the construction of our instruction finetuning dataset and the corresponding experiments. We evaluate the proposed method among mainstream LLMs and present the results in the subsequent subsections.
4.1 Data Construction
Following the proposed two-stage framework, we synthesize two new instruction datasets for LLM finetune. Besides, we construct a prefix uncertainty-recognition evaluation dataset to better evaluate the LLMs’ recognition of not-knowing.
Uncertainty-Recognition Benchmark To comprehensively evaluate the model’s cognitive ability to recognize knowledge gaps, we construct a testbed using the ASQA Stelmakh et al. (2022) dataset, which consists of ambiguous questions. Each question is divided into multiple sub-questions with their corresponding contexts. For example, one pair may discuss the discovery of the photoelectric effect in 1887, while another may cover the theoretical development in 1905. To generate the unknown questions, we shuffle these pairs, reassigning the questions to different but related contexts. As a result, there are two advancements of our benchmark differentiating from other works: First, the context is closely relevant to the query with partially date or object mismatched, for challenging the model’s ability to handle uncertainty; Second, context is crucial but insufficient to the query. Such samples are rare and valuable, as ASQA is the only dataset we have found that could generate sufficient samples that satisfy the requirement. During the evaluation, two types of designed evaluation templates are available: the QA uncertainty recognition template and the postfix uncertainty recognition template. For the QA uncertainty recognition, we instruct the model to either answer the question or reject it with ’Not Provided’. We generate 3,320 samples for this type of template. For postfix uncertainty recognition, we instruct the LLMs to identify the sufficiency after generation. We include 646 test samples for this template. Here is a demonstration of the used prompt templates:
  • QA Uncertainty-Recognition: If the context is not sufficient to answer the question, please answer it with ’Not Provided’
  • Postfix Uncertainty-Recognition: You must append either ’<Sufficient>’ or ’<Insufficient>’ after your answer.
Uncertainty-Sensitive Tuning Dataset In uncertainty-sensitive tuning, two distinct instruction datasets are used for two stages separately: the binary uncertainty-recognition tuning dataset and the prompt-sensitive tuning dataset. For the uncertainty-recognition tuning, we sample 3,320 examples from the ASQA Stelmakh et al. (2022) dataset with the ground truth hidden, ensuring no overlap with the evaluation data. To protect our valuable benchmark, the prompt-sensitive tuning is based on HotpotQA Yang et al. (2018), a dataset designed for multi-hop reasoning QA. To adapt it with additional causal instructions, we generate and manually select the most effective 28 causal instructions. Then, we integrate these instructions with 300 randomly selected samples from HotpotQA. Finally, all the generated datasets are used for fine-tuning. Here is a demonstration of the additional prompt in uncertainty-recognition tuning:
  • Uncertainty-Recognition Tuning: You must only answer either ’Sufficient’ or ’Insufficient’ without any other output
QA Uncertainty-Recognition
Postfix
Uncertainty-Recognition
Category Model $Acc_{known}$ $Acc_{unknown}$ $F1$ $Acc_{known}$ $Acc_{unknown}$ $F1$
Benchmark GPT-4 Turbo 79.6 83.6 81.6 88.9 78.3 83.3
GPT-3.5 Turbo 82.1 51.8 63.5 97.0 33.4 49.7
Vicuna-7B v1.5 74.6 43.8 55.2 93.5 14.3 24.8
Self-RAG-7B 67.9 48.1 56.3 46.0 74.9 57.0
Llama2 Llama2-7B-Chat(baseline) 79.3 58.3 67.2 85.2 29.5 43.9
Uncertainty-Recognition Tuned 52.4 84.4 64.6 81.3 84.0 82.6
Uncertainty & HotpotQA Tuned 77.0 20.9 32.8 87.1 66.7 75.5
Uncertainty-Sensitivity Tuned 79.7 93.0 85.8 88.0 66.0 75.4
Mistral Mistral-7B-Instruct-v0.2(baseline) 85.1 63.0 72.4 82.8 43.1 56.7
Uncertainty-Recognition Tuned 77.5 75.8 76.6 86.1 81.9 84.0
Uncertainty & HotpotQA Tuned 87.1 52.4 65.5 80.7 75.1 77.8
Uncertainty-Sensitivity Tuned 87.3 75.3 80.9 82.5 82.2 82.4
Gemma Gemma-2-9B-Instruct(baseline) 86.1 74.1 73.5 86.3 57.6 69.1
Uncertainty-Recognition Tuned 76.1 86.2 80.8 93.4 76.2 83.9
Uncertainty & HotpotQA Tuned 91.3 20.8 33.9 99.4 58.7 73.8
Uncertainty-Sensitivity Tuned 87.6 81.2 84.3 96.1 55.1 70.1
Table 2: Results (in %) for mainstream and our finetuned LLMs on QA uncertainty-recognition assessment and postfix uncertainty-recognition assessment. The overall best results are highlighted in<textbfbold. The results obviously higher and lower than the baseline are highlighted in green and orange, respectively.
4.2 Experiment Setting
Training Details. We experiment our uncertainty-sensitive tuning approach on mainstream open-sourced LLMs, including Llama2-7B-Chat Touvron et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), and Gemma-2-9B-Instruct Team et al. (2024). We also test GPT-4 Turbo OpenAI (2023), GPT-3.5 Turbo OpenAI (2023), Vicuna-7B v1.5 Zheng et al. (2024a) and Self-RAG-7B Asai et al. (2023) on our benchmark. Our finetuning bases on a single RTX3090 GPU by LLaMA-Factory Zheng et al. (2024b), with Lora Hu et al. (2021) in a rank of 8, with a batch size of 4 and a learning rate of 5e-5. This research integrates the instruction-based and attributed prompt which is proved to be effective for mitigating hallucination Zhou et al. (2023), as provided in Appx. A.
Evaluation Metric. We use $Acc_{known}$ for representing the accuracy of questions with specific answers, and $Acc_{unknown}$ for unknown questions.
4.3 Results
Table 2 presents the results of our benchmark. The evaluation of QA uncertainty-recognition reveals that most mainstream LLMs struggle to accurately identify unknown questions, with only around 60% accuracy. However, GPT-4 and Gemma-2 are notable exceptions with accuracies of 83.6% and 74.1%, respectively. Mistral and Llama-2 rank highest among the remaining models, even surpassing GPT-3.5, despite the latter having more parameters. But there is still a substantial performance gap between the GPT-4 and the other models. The ongoing experiments aim to investigate the reasons behind this disparity. The postfix uncertainty-recognition assessment results, shown in the right column of Table 2, indicate a strong positive correlation between the two evaluation templates. The analysis also highlights that different models respond differently to insufficient queries. Models fine-tuned on dialogue tasks tend to overly rely on and trust the given information. Self-RAG, which is fine-tuned for QA tasks with unknown questions, demonstrates a strong ability to identify the questions it cannot answer. However, it remains difficult for the model to admit the lack of knowledge.
Our uncertainty-recognition tuned models perform comparably to GPT-4 in tests of uncertainty recognition, and show an improvement up to 25.9% in responding to unknown questions compared to the baselines. However, despite outperforming GPT-4 in terms of unknown questions ($Acc_{unknown}$), this increased awareness of uncertainty leads to a decrease in original QA capabilities. Models are overly sensitive to the diverse ways in which users might phrase similar questions. Further fine-tuning on the HotpotQA resulted in the model being less reliable in following instructions and sometimes disregarding its knowledge limitations, despite the advancements in postfix uncertainty-recognition indicates the knowledge is maintained. Among the models tested, the uncertainty-sensitivity tuned Llama2 ranks the highest, achieving an F1 score of 85.8, which is 4.2 points higher than GPT-4 and 18.6 points above the baseline (as shown in Figure 2). This impressive performance is due to the model’s ideal balance between QA and rejecting answering. It achieves a remarkable 93.0% accuracy on unknown questions, the highest among all mainstream LLMs, while maintaining a 79.7% accuracy on known questions. Gemma-2 and Mistral also achieve 10.8% and 8.5% improvements in overall performance, demonstrating that our uncertainty-sensitivity tuning approach allows models to effectively integrate new information without compromising the knowledge acquired from uncertainty-recognition fine-tuning.
4.4 Ablation Study
4.4.1 Experiment Setting
To further quantify the effects of uncertainty-recognition tuning, HotpotQA tuning and prompt-sensitive tuning, three additional variants have been built on Llama2-7B-Chat for ablation experiments.
  • The HotpotQA fine-tuning utilizes 300 samples with traditional question answering prompts provided in Appx. A.1.
  • The prompt-sensitive tuning bases on the 300 samples mentioned in Section 3.4.
  • The Uncertainty & HotpotQA tuned model further fine-tunes the uncertainty-recognition tuned model on HotpotQA dataset.
4.4.2 Result Analysis
Table 2 demonstrates that the reasons for LLMs’ poor performance in rejecting the unknown questions come from two aspects: weak uncertainty-recognition capacity and the prompt-sensitivity reduction problem.
Weak Uncertainty-Recognition Capacity: Our ablation study, as shown in Table 3, illustrates that enhancing the model’s recognize-uncertainty ability significantly boosts performance on unknown questions. Specifically, the uncertainty-recognition tuning resulted in a marked 16.9% increase in ${Acc}_{unknown}$. The problem of weak uncertainty-recognition capacity in LLMs comes from their fundamental training process to generate specific responses. In scenarios where LLMs encounter questions outside their training distribution or scope of knowledge, they often fail to adequately recognize uncertainty. This leads to inappropriate or misleading answers.
QA Uncertainty-Recognition
Model $Acc_{known}$ $Acc_{unknown}$ $F1$
Llama2-7B-Chat (baseline) 79.3 58.3 67.2
Uncertainty-Recognition Tuned 52.4 84.4 64.6
HotpotQA Tuned 77.5 58.3 66.5
Prompt-Sensitive Tuned 84.8 59.0 69.6
Uncertainty & HotpotQA 77.0 20.9 32.8
Uncertainty-Sensitive Tuned 79.7 93.0 85.8
Table 3: Results of ablation. The overall best results are highlighted in bold. The results obviously higher and lower than the baseline are highlighted in green and orange, respectively.
Figure 3: Comparison of different models in ablation study. A higher score means better performance.
Prompt-sensitivity reduction problem: The reduction in known questions is another critical issue highlighted by the ablation study. As models are further tuned on specific tasks, such as QA or uncertainty recognition, they can become less responsive to variations in the input prompt, a phenomenon we term "prompt-sensitivity reduction." This reduced sensitivity can impair the model’s ability to adapt to the slight but significant differences in how questions are phrased or presented, leading to what we describe as cognitive inaccessibility. From Table 3, it is evident that although tuning on HotpotQA attempts to mitigate the negative effects of over-tuning for uncertainty-recognition, it does not fully address the prompt-sensitivity issue. The model even ignores the constraints in prompts and continues to hallucinate answers for unknown questions. Thus, the model fails to fully leverage the context of the prompt, resulting in inaccurate or relevant answers. With our prompt-sensitive tuning, we can successfully recover the performance of other questions. Even if we do not fine-tune the model for the unknown task, our prompt-sensitive tuning can boost the performance of the task with specific answers, which is the highest accuracy in the table.
4.5 Comparisons with Post-Generation Methods
Cole et al. (2023) and Kadavath et al. (2022) propose using sampling and LM-validation for post-generation uncertainty detection. Here we compare their approaches with uncertainty-sensitive tuning. For sampling method, we sample three times from the Llama2-7B-Chat with 0.6 temperature, and adopt the most repeated answers. In LM-Validation we allow the model to additionally edit the generated answers. As recorded in Table 4, due to the weak uncertainty-recognition capacity, models struggle to identify the unknown questions.
Model $Acc_{pos}$ $Acc_{neg}$ $F1$
Llama2-7B-Chat (baseline) 79.3 58.3 67.2
Multi-Sampling 79.7 66.5 72.5
LM-Validation 82.5 53.8 65.1
Uncertainty-Sensitivity Tuned 79.7 93.0 85.8
Table 4: Comparison results with the post-generation methods on QA uncertainty-recognition benchmark.
5 Conclusion
In this paper, we explore a common issue in large language models (LLMs) where insufficient context in prompts leads to the generation of plausible yet incorrect responses. Firstly, our research identifies that LLMs often fail to recognize when they lack adequate information to accurately respond. We primarily attribute this problem to the models’ weak uncertainty-recognition capacity and overconfidence in their generated potential answers. Second, training on the questions with specific answers significantly decreases the performance of questions without answers. To counter these issues, we introduce a novel two-stage training framework termed "uncertainty-sensitive tuning". In the first stage, we guide the LLM to identify unknown questions. In the second stage, we recover the decreased QA performance by incorporating designed causal instructions. This approach significantly enhances the model’s reliability, reducing the occurrence of hallucinations. Our approach differs from previous works in that it focuses more directly on refining the response behavior through targeted fine-tuning, thereby enhancing its performance in practice. By open-sourcing this framework and dataset, we believe that it will highlight a new direction for the automatic synthesis instruction dataset, which is not only focused on the diversity of the data but also the reduction of hallucinations during training.
Limitations
In this work, we have two promising directions to refine in the future. The first limitation is that when tuned with the new datasets, which contain a limited number of unknown questions, the LLM experiences a long-tail problem. Therefore, to effectively tailor our proposed two-stage, uncertainty-sensitive method to specific datasets, it requires further tuning. Second, we face a limitation in that we did not probe or analyze the parametric knowledge learned by Llama2 during its initial pre-training phase. Our fine-tuning dataset may overlap with Llama2’s pre-training data. To address these challenges, our future work aims to explore existing methods that measure model uncertainty by monitoring internal parameters Lu et al. (2023). By quantifying the model’s uncertainty on different inputs, we can gain insights into its knowledge gaps and long-tail weaknesses. This knowledge could then lead to more targeted fine-tuning strategies or engineering methods that make the fine-tuned LLM work better on different types of queries.
Ethics Statement
Our benchmark and datasets are constructed from public datasets, and uncertainty-sensitive tuning dataset further introduces GPT-4 to refine the dataset, thus inherent biases may exist. However, the techniques utilized in this work do not introduce additional biases beyond what the data has presented.
Appendix A Instructions
A.1 Question Answering
Question Answering Task: You need to do the Question Answering for the following query. I will give a question and several contexts. Based on the given contexts, give an answer to the question. Your answer must not using any additional knowledge that is not mentioned in the contexts. If the context is not sufficient to answer the question, please answer it with ’Not Provided’ QUERY: $q$ CONTEXT: $c$ ANSWER:
A.2 Uncertainty-Recognition
Cognition Assessment Task: You need to do the Cognition Assessment for the following query. I will give a query and a related context about the query. Your task is to judge whether the context is sufficient to answer the query. Assessment: You must append either ’<Sufficient>’ or ’<Insufficient>’ after your answer. Finetuning: You must only answer either ’Sufficient’ or ’Insufficient’ without any other output. Here is the example. QUERY: What happened to Jay when he got old? CONTEXT: Jay Chou was the most famous singer in China when he was young, releasing many nostalgic albums and songs that are memorable to middle-aged people today. ANSWER: Assessment: Jay Chou was the most famous singer in China.<Insufficient> Finetuning: Insufficient Here is the provided information that you need to accomplish follow the provided example: QUERY: $q$ CONTEXT: $c$ ANSWER:
A.3 Additional Casual Instruction Generation
Your task is to provide various instructions for the questions answering task. The questions answering task provides a context and a query. e.g. "Context: XXX Query: XXX Answer:". And your task is to add some specific requirement to the answer. e.g. "The answer must be all in upper case", "There should be no punctuation in the answer". The added instruction should be general to the query. You should generate hundreds of such instructions.
A.4 Prompt-Sensitive Finetuning
You should check whether your answer aligned the requirement by generating a Checking part, checking each sentence of the above instruction, with either <fulfilled> or <not fulfilled> mark behind the sentence, indicating whether the requirement is fulfilled or not. If there is <not fulfilled> mark behind the sentence, you must modify your answer again to fulfill the requirement, by appending a new ANSWER and CHECKING part. Here is an example for this task: e.g. Question Answering Task Requirements: You need to do the Task Prompt for the following query and context. Ensure the response is written in the past tense. QUESTION: Who is Jack Chen? CONTEXTS: People saying that Jack Chen is a famous singer in China. ANSWER: Jack Chen is a famous singer in China. CHECKING: Question Answering Task: You need to do the Task Prompt for the following query and context.<fulfilled>Ensure the response is written in the past tense.<not fulfilled> ANSWER: Jack Chen is a famous singer in China. CHECKING: Question Answering Task: You need to do the Task Prompt for the following query and context.<fulfilled>Ensure the response is written in the past tense.<fulfilled> Here is the information of your task: {Question Answering Instruction}
A.5 Trustworthy Question Answering for Benchmark
Trustworthy Question Answering Task: You need to utilize the ability learnt during both the Question Answering Task and Cognition Assessment Task. And only provide the answers which are sufficiently supported by the context, otherwise provide ’Not Provided’ I will give a question and several context texts about the question. Based on the given contexts, give an answer to the question. Your answer must not using any additional knowledge that is not mentioned in the given contexts. If the context is not sufficient to answer the question, please answer it with ’Not Provided’ QUERY: $q$ CONTEXT: $c$ ANSWER:
Appendix B Example of Proposed Benchmark
Question: Who discovered and developed an explanation for the
photoelectric effect in 1887?
Positive Context: This is a passage about Photoelectric effect:
Light, and especially ultra-violet light, discharges negatively electrified
bodies with the production of rays of the same nature as cathode rays.
Under certain circumstances it can directly ionize gases. The first of
these phenomena was discovered by Heinrich Hertz and Wilhelm
Hallwachs in 1887. The second was announced first by Philipp Lenard in 1900.
Negative Context: This is a passage about Photoelectric effect:
In 1905, Einstein proposed an explanation of the photoelectric effect using a
concept first put forward by Max Planck that light waves consist of tiny
bundles or packets of energy known as photons or quanta.